# GraphReader: Building Graph-based Agent to Enhance Long-Context Abilities of Large Language Models

### v1

@@ Abstract
Shilong Li∗1, Yancheng He∗1, Hangyu Guo∗1, Xingyuan Bu∗†‡1, Ge Bai1, Jie Liu2,3, Jiaheng Liu1, Xingwei Qu4, Yangguang Li3, Wanli Ouyang2,3, Wenbo Su1, Bo Zheng1 1Alibaba Group 2The Chinese University of Hong Kong 3Shanghai AI Laboratory 4University of Manchester {zhuli.lsl, buxingyuan.bxy}@taobao.com
@@ 1 Introduction
Large language models (LLMs) have made great progress on natural language understanding and generation (Zhao et al., 2023). However, transformer-based LLMs still struggle in handling long contexts due to the limitation of context window and memory usage.
Current techniques for solving the long-context tasks of LLMs can be divided into two perspectives: 1) Model-level, which includes finetuning with modified positional embeddings Chen et al. (2023b); Zhu et al. (2023); Peng et al. (2023); Ding et al. (2024), and applying transformer variants with modified attention mechanisms (Dai et al., 2019; Munkhdalai et al., 2024; Gu and Dao, 2023); 2) Agent-level, i.e., employing retrieval-augmented LLM or agent to process long contexts with a limited context window LLM Nakano et al. (2021); Lee et al. (2024).
However, model-level methods typically train LLMs with target length texts, posing challenges in constructing training datasets and incurring high training costs Zhu et al. (2023). Additionally, long-context LLMs optimized with these methods tend to overlook crucial details in long contexts, known as “lost in the middle” Liu et al. (2024), limiting their ability to address complex tasks, such as multi-hop questions. Agent-level approaches transform input text into a tree (Chen et al., 2023a) or paginated pages (Lee et al., 2024), failing to capture multi-hop and long-range dependencies, thus limiting their effectiveness on very long contexts, as shown in Figure 1.
@@ 2 Related Work
Recent efforts (Chen et al., 2023b; Ding et al., 2024; Peng et al., 2023) have focused on positional interpolation (PI) to enhance long-context capabilities. However, these methods require training on full-length texts, leading to significant increases in data and training costs Chen et al. (2023c); Fu et al. (2024); Bai et al. (2024b). Thus, PoSE Zhu et al. (2023) and SkipAlign Wu et al. (2024a) investigate data skip strategy, but tend to neglect detailed information in long texts Liu et al. (2024); Bai et al. (2024a); Wu et al. (2024b). Furthermore, despite how extensively the context window is expanded, it remains constrained by a predefined fixed length. To address these limitations, transformer variants with modified attention mechanisms have been proposed (Dai et al., 2019; Gu and Dao, 2023; Munkhdalai et al., 2024). However, these models are prone to losing earlier information.
@@ 3.1 Preliminary
GraphReader is built on a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E}\right\}$, where each node $v_{i}\in\mathcal{V}$ contains a key element $k_{i}$ and a set of summarized content, namely atomic facts $\mathcal{A}_{i}$. In other words, $v_{i}=\left\{k_{i},\mathcal{A}_{i}\right\}$. And each edge $e_{ij}\in\mathcal{E}$ represents the relationship between nodes $v_{i}$ and $v_{j}$. This graph structure enables GraphReader to capture global information from the input document $D$ within a limited context window, allowing it to decide whether to explore the current node in detail or jump to a neighboring node. During graph exploration, GraphReader collects supporting facts and terminates the exploration once sufficient information has been gathered to answer the question. As illustrated in Figure 2, the entire process of GraphReader consists of the following three phases: graph construction, graph exploration, and answer reasoning. The prompts utilized in these three stages are detailed in Appendix A, and a detailed example of our process can be found in Appendix H.
@@ 4.1 Experimental Settings
We compare our approach with the following baselines: retrieval augmented generation (RAG), long-context LLM, and agent-based methods. (1) RAG: We choose Okapi BM25 (Robertson and Zaragoza, 2009) or OpenAI API embedding model Ada-002 https://platform.openai.com/docs/guides/embeddings/embedding-models to retrieve the chunks most relevant to the question and employ GPT-4-128k (gpt-4-1106-preview) to read retrieved chunks and answer the question. (2) Long-context LLM: We select GPT-4-128k for directly reading full text when the text content fits within the input window, or for segmenting the text into chunks for sequential reading. (3) Agent-based Method: We select ReadAgent (Lee et al., 2024), which employs an agent-based system for the execution of retrieval and reading processes for long-context QA. The detailed description of these methods is provided in Appendix D.
@@ 4.4 Further Analysis
To evaluate the inference cost of our approach, we compare the average token consumption of ReadAgent and GraphReader for each question. As shown in Table 5, GraphReader uses only 1.08 times more tokens than ReadAgent but achieves over twice the performance improvement, demonstrating its superiority. Additionally, after the first round of document preprocessing, token consumption in subsequent exploration is significantly reduced and much lower than the original dataset’s token count.
@@
To evaluate our method’s advantages in key information recall, we utilize GPT-4 to assess the recall of supporting facts on the HotpotWikiQA-mixup dataset. As shown in Figure 5, our model consistently outperforms other baseline methods, regardless of the input length. As context length increases from 16k to 256k, recall of supporting facts declines across all methods. However, GraphReader maintains around 60% recall at 256k context length, in contrast to the significant degradation in ReadAgent. This demonstrates GraphReader’s scalability and effectiveness in processing long contexts. Further details and evaluation prompts can be found in Appendix E.
To further demonstrate the recall rate of GraphReader at different granularities, we calculate the recall rate of Supporting Facts and Sample granularity respectively using the same method, detailed in the Appendix E. The granularity of supporting facts refers to the recall rate of all supporting facts across the entire dataset. As for sample granularity, a sample is considered to be recalled only if all of its supporting facts are recalled. As shown in the Tabel 6, the recall for the final notebook is slightly higher than the recall of atomic facts, which indicates that our method is capable of extracting more valid information from chunks during the exploration, indirectly reflecting its intelligence and effectiveness in exploration.
@@ Appendix D Baseline Methods
RAG is a commonly used approach for addressing long-text problems. In this work, we compare the traditional RAG method, including retrieval methods based on Okapi BM25 (Robertson and Zaragoza, 2009) and the OpenAI API embedding model (text-embedding-ada-002). Specifically, we first split the text into chunks in the same method as GraphReader, then use the aforementioned methods to calculate the relevance scores between the question and these chunks, and finally input the top-n chunks with the highest relevance scores together with the question for the LLM to answer.
To ensure a fair comparison, we control the input window to 4k in the experiments. Specifically, in order to fill the input window as much as possible, we set the maximum length of the chunk to 38k when selecting the top-1 chunk for answering; when opting for the top-3 chunks, we set the maximum length of each chunk to 1k. The specific prompt can be found in Figure 18.
@@
Appendix E Evaluation Recall for Supporting Facts
@@
Appendix F The Analysis of Function Calls
@@
Appendix G Statistics of Graph
@@
Appendix H GraphReader Example
@@
Figure 20: GraphReader Example(Question and Annotations). We provide an example question with its answer, along with the supporting passages for this question. This is a typical 3-hop question where we need to gather information and reason step-by-step to arrive at the answer.

### v2

@@ Abstract
Shilong Li∗1, Yancheng He∗1, Hangyu Guo∗1, Xingyuan Bu∗†‡1, Ge Bai1, Jie Liu2,3, Jiaheng Liu1, Xingwei Qu4, Yangguang Li3, Wanli Ouyang2,3, Wenbo Su1, Bo Zheng1 1Alibaba Group 2The Chinese University of Hong Kong 3Shanghai AI Laboratory 4University of Manchester zhuli.lsl@taobao.com, xingyuanbu@gmail.com
@@ 1 Introduction
Large language models (LLMs) have made great progress on natural language understanding and generation (Zhao et al., 2023; Liu et al., 2024a; Feng et al., 2022; Peng et al., 2020; Xv et al., 2022; Peng et al., 2023b; Bu et al., 2021). However, transformer-based LLMs still struggle in handling long contexts due to the limitation of context window and memory usage.
Current techniques for solving the long-context tasks of LLMs can be divided into two perspectives: 1) Model-level, which includes finetuning with modified positional embeddings Chen et al. (2023b); Zhu et al. (2023); Peng et al. (2023a); Ding et al. (2024), and applying transformer variants with modified attention mechanisms (Dai et al., 2019; Munkhdalai et al., 2024; Gu and Dao, 2023); 2) Agent-level, i.e., employing retrieval-augmented LLM or agent to process long contexts with a limited context window LLM Nakano et al. (2021); Lee et al. (2024).
However, model-level methods typically train LLMs with target length texts, posing challenges in constructing training datasets and incurring high training costs Zhu et al. (2023). Additionally, long-context LLMs optimized with these methods tend to overlook crucial details in long contexts, known as “lost in the middle” Liu et al. (2024b), limiting their ability to address complex tasks, such as multi-hop questions. Agent-level approaches transform input text into a tree (Chen et al., 2023a) or paginated pages (Lee et al., 2024), failing to capture multi-hop and long-range dependencies, thus limiting their effectiveness on very long contexts, as shown in Figure 1.
@@ 2 Related Work
Recent efforts (Chen et al., 2023b; Ding et al., 2024; Peng et al., 2023a) have focused on positional interpolation (PI) to enhance long-context capabilities. However, these methods require training on full-length texts, leading to significant increases in data and training costs Chen et al. (2023c); Fu et al. (2024); Bai et al. (2024b). Thus, PoSE Zhu et al. (2023) and SkipAlign Wu et al. (2024a) investigate data skip strategy, but tend to neglect detailed information in long texts Liu et al. (2024b); Bai et al. (2024a); Wu et al. (2024b). Furthermore, despite how extensively the context window is expanded, it remains constrained by a predefined fixed length. To address these limitations, transformer variants with modified attention mechanisms have been proposed (Dai et al., 2019; Gu and Dao, 2023; Munkhdalai et al., 2024). However, these models are prone to losing earlier information.
@@ 3.1 Preliminary
GraphReader is built on a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E}\right\}$, where each node $v_{i}\in\mathcal{V}$ contains a key element $k_{i}$ and a set of summarized content, namely atomic facts $\mathcal{A}_{i}$. In other words, $v_{i}=\left\{k_{i},\mathcal{A}_{i}\right\}$. And each edge $e_{ij}\in\mathcal{E}$ represents the relationship between nodes $v_{i}$ and $v_{j}$. This graph structure enables GraphReader to capture global information from the input document $D$ within a limited context window, allowing it to decide whether to explore the current node in detail or jump to a neighboring node. During graph exploration, GraphReader collects supporting facts and terminates the exploration once sufficient information has been gathered to answer the question. As illustrated in Figure 2, the entire process of GraphReader consists of the following three phases: graph construction, graph exploration, and answer reasoning. The prompts utilized in these three stages are detailed in Appendix A, and a detailed example of our process can be found in Appendix I.
@@ 4.1 Experimental Settings
We compare our approach with the following baselines: retrieval augmented generation (RAG), long-context LLM, and agent-based methods. (1) RAG: We choose Okapi BM25 (Robertson and Zaragoza, 2009) or OpenAI API embedding model Ada-002 https://platform.openai.com/docs/guides/embeddings/embedding-models to retrieve the chunks most relevant to the question and employ GPT-4-128k (gpt-4-1106-preview) to read retrieved chunks and answer the question. In addition to traditional RAG methods, we also compared GraphRAG (Edge et al., 2024) and LongRAG (Jiang et al., 2024), which utilize LLM to enhance RAG ability. (2) Long-context LLM: We select GPT-4-128k for directly reading full text when the text content fits within the input window, or for segmenting the text into chunks for sequential reading. (3) Agent-based Method: We select ReadAgent (Lee et al., 2024) and PEARL (Sun et al., 2024), which employ an agent-based system for the execution of retrieval and reading processes for long-context QA. The detailed description of these methods is provided in Appendix D.
@@ 4.2 Main Results
For additional results on benchmarks relevant to real-world scenarios, please refer to the appendix E.
@@ 4.4 Further Analysis
To assess the inference cost of our approach, we compare the average token consumption of ReadAgent and GraphReader for individual questions. As shown in Table 5, GraphReader uses only 1.08 times more tokens than ReadAgent (52.8k / 48.7k), yet achieves more than double the performance improvement, demonstrating its superiority. More importantly, our method has significant advantages in single-document multiple-query scenarios, where only one graph needs to be constructed. Subsequent QA can be performed on this graph, thereby reducing the overall token consumption.
@@
To evaluate our method’s advantages in key information recall, we utilize GPT-4 to assess the recall of supporting facts on the HotpotWikiQA-mixup dataset. As shown in Figure 5, our model consistently outperforms other baseline methods, regardless of the input length. As context length increases from 16k to 256k, recall of supporting facts declines across all methods. However, GraphReader maintains around 60% recall at 256k context length, in contrast to the significant degradation in ReadAgent. This demonstrates GraphReader’s scalability and effectiveness in processing long contexts. Further details and evaluation prompts can be found in Appendix F.
To further demonstrate the recall rate of GraphReader at different granularities, we calculate the recall rate of Supporting Facts and Sample granularity respectively using the same method, detailed in the Appendix F. The granularity of supporting facts refers to the recall rate of all supporting facts across the entire dataset. As for sample granularity, a sample is considered to be recalled only if all of its supporting facts are recalled. As shown in the Tabel 6, the recall for the final notebook is slightly higher than the recall of atomic facts, which indicates that our method is capable of extracting more valid information from chunks during the exploration, indirectly reflecting its intelligence and effectiveness in exploration.
@@ Appendix C Dataset
Real-World Datasets
QuALITY (Pang et al., 2022) is a long-text multiple-choice question-answering dataset, with questions crafted by contributors who are familiar with the complete passages, making it more representative of real-world QA scenarios. We handle it as straightforward QA problems. Natural Questions (Kwiatkowski et al., 2019) includes real anonymous aggregated queries from Google along with corresponding Wikipedia pages, providing another excellent resource for authentic long-text QA situations.
@@ Appendix D Baseline Methods
RAG is a commonly used approach for addressing long-text problems. In this work, we compare the traditional RAG method, including retrieval methods based on Okapi BM25 (Robertson and Zaragoza, 2009) and the OpenAI API embedding model (text-embedding-ada-002). Specifically, we first split the text into chunks in the same method as GraphReader, then use the aforementioned methods to calculate the relevance scores between the question and these chunks, and finally input the top-n chunks with the highest relevance scores together with the question for the LLM to answer. To ensure a fair comparison, we control the input window to 4k in the experiments. Specifically, in order to fill the input window as much as possible, we set the maximum length of the chunk to 38k when selecting the top-1 chunk for answering; when opting for the top-3 chunks, we set the maximum length of each chunk to 1k. The specific prompt can be found in Figure 18.
In addition to traditional RAG methods, we also compared GraphRAG (Edge et al., 2024) and LongRAG (Jiang et al., 2024). GraphRAG utilizes LLMs to construct a graph-based text index in two distinct stages. The first stage involves extracting an entity knowledge graph from the source documents, while the second stage focuses on generating summaries for groups of entities. When a question is posed, each summary provides partial information, which is then combined into the final answer for the user. LongRAG introduces a “long retriever” and a “long reader”, allowing the entire corpus to be processed into larger-sized units, which reduces the number of units needed during retrieval and alleviates the burden on the retriever.
@@
We also compared PEARL (Sun et al., 2024), a prompting framework to enhance reasoning capabilities for long documents. PEARL is structured into three stages: action mining, plan formulation, and plan execution. It decomposes complex questions into actionable steps and utilizes LLMs for zero-shot or few-shot prompting execution.
Appendix E Additional Experimental Results
Method Input QuALITY Natural Question
Window LR-1 LR-2 EM $F_{1}$ LR-1 LR-2 EM $F_{1}$
GPT-4-128k $128k$ 45.7 60.3 2.7 9.9 75.0 81.0 41.0 57.2
Pearl $128k$ 52.3 72.7 3.0 9.6 74.0 79.0 38.0 56.8
LongRAG $128k$ 52.3 67.0 3.7 11.6 77.7 83.0 47.3 60.0
GraphRAG $128k$ 46.0 68.7 2.0 6.1 67.0 77.0 47.0 53.2
GraphReader $4k$ 57.3 82.3 4.3 14.3 79.0 84.7 48.3 62.1
Table 7: Performance (%) comparison of different baselines on two additional datasets. The best performance and the second-best performance are denoted in bold and underlined fonts, respectively.
Table 7 presents additional experimental results for two datasets, QuALITY and Natural Questions, both of which are highly relevant to real-world question-answering scenarios. The results indicate that our method significantly outperforms other baseline models in real-world scenarios.
Appendix F Evaluation Recall for Supporting Facts
@@
Appendix G The Analysis of Function Calls
@@
Appendix H Statistics of Graph
@@
Appendix I GraphReader Example
@@
Figure 20: GraphReader Example(Question and Annotations). We provide an example question with its answer, along with the supporting passages for this question. This is a typical 3-hop question where we need to gather information and reason step-by-step to arrive at the answer. Details about this example are available at this link.